# KITchen: A Real-World Benchmark and Dataset for 6D Object Pose Estimation in Kitchen Environments

### v1

@@ Abstract
Despite the recent progress on 6D object pose estimation methods for robotic grasping, a substantial performance gap persists between the capabilities of these methods on existing datasets and their efficacy in real-world mobile manipulation tasks, particularly when robots rely solely on their monocular egocentric field of view (FOV). Existing real-world datasets primarily focus on table-top grasping scenarios, where a robotic arm is placed in a fixed position and the objects are centralized within the FOV of fixed external camera(s). Assessing performance on such datasets may not accurately reflect the challenges encountered in everyday mobile manipulation tasks within kitchen environments such as retrieving objects from higher shelves, sinks, dishwashers, ovens, refrigerators, or microwaves. To address this gap, we present Kitchen, a novel benchmark designed specifically for estimating the 6D poses of objects located in diverse positions within kitchen settings. For this purpose, we recorded a comprehensive dataset comprising around 205k real-world RGBD images for 111 kitchen objects captured in two distinct kitchens, utilizing one humanoid robot with its egocentric perspectives. Subsequently, we developed a semi-automated annotation pipeline, to streamline the labeling process of such datasets, resulting in the generation of 2D object labels, 2D object segmentation masks, and 6D object poses with minimized human effort. The benchmark, the dataset, and the annotation pipeline are available at https://kitchen-dataset.github.io/KITchen.
@@ I Introduction
To overcome the limitations of the current 6D pose estimation methods, we introduce KITchen, the first-of-its-kind large-scale real-world dataset recorded using the humanoid robot ARMAR-6 [3], which has adjustable height and roll-yaw neck, in 2 different kitchen environments covering 111 kitchen objects from the humanoid robots’ egocentric perspective to cover the objects in the challenging kitchens’ locations as shown in Fig. 1. KITchen offers 2D bounding boxes, object segmentation, and 6D poses annotated with a semi-automated annotation pipeline to minimize the need for manual labeling.
The main contributions of our work are: (i) we introduce a large real-world annotated RGBD dataset for 111 objects with their 2D bounding boxes, segmentation masks, and 6D poses. (ii) we propose a semi-automated annotation pipeline to annotate the objects in the dataset to facilitate the creation of more real-world datasets and make it publicly available to support other robotics groups to create such large-scale datasets. (iii) we introduce a new benchmark and competition, where the focus is to solve the object 6D pose estimation problem depending solely on the monocular FOV of robots and limiting the submissions to approaches that offer at least $5fps$ to encourage further work on this problem while taking into consideration real-time applicability.
@@ II Related Work
Dataset Objects Images Annotated Objects/Image Multi-object Multi-instance Robot’s FOV
@@
TABLE I: Overview of available datasets for instance-level 6D pose estimation, highlighting key metrics including the number of covered objects in the dataset, total images count, number of annotated objects per image, presence of multi-object setups, availability of multiple instances of same objects, and whether the dataset was captured using a mobile robot’s field of view.
@@ II-A Objects Datasets
KIT object models database [28] was originally introduced in 2012 and offers 3D CAD models for more than 100 diverse objects, the majority of which are kitchen-related groceries. However, it only offers very few images for each object, which makes it hard to use this dataset for 6D pose estimation with the current state-of-the-art (SOTA) data-driven 6D pose estimation approaches. KIT bimanual manipulation dataset [30] provides rich data for learning models of bimanual manipulation tasks from human demonstrations. It includes accurate whole-body motion data, hand configurations, and 6D object poses captured using various sensors. The dataset features 12 bimanual actions for 21 kitchen-related objects.
@@ III-A Dataset’s Objects
We aim to create a large-scale real-world dataset that covers objects that are commonly used inside the kitchen environments. Although some of the existing object datasets already offer objects that are commonly used in kitchens, they lack enough diverse RGBD annotated images to train on [49] or no annotated RGBD at all [7, 30, 28, 34]. Therefore, we decided to reuse the already available kitchen-related objects from these datasets and provide a large real-world RGBD annotated dataset for them to facilitate research on 6D pose estimation for kitchen objects. These objects vary from toy vegetables and fruits from [7] to kitchen tools such as knives, spoons, cups, mugs, bowls, cutting board, egg whisk, frying pan, plate, etc from [7, 49, 30, 28, 34] to kitchen groceries objects from [7, 28].
@@ III-B Dataset Recording
Figure 3: Diverse robot torso positions. The images display heights of 145cm, 177cm, and 185cm from left to right, illustrating the varied perspectives captured in the datasets and the different placements of objects relative to the robot’s field of view.
We recorded the dataset using our humanoid robot ARMAR-6 [2, 4] inside two distinct kitchen environments as seen in Fig. 2 the first kitchen, referred to as the Main Kitchen, includes typical kitchen appliances such as a fridge, counter with drawers, table, sink, microwave, dishwasher, and oven. The second kitchen, named Mobile Kitchen, features a counter with drawers, sink, dishwasher, fridge, and three tables. To enhance diversity, we utilized four different table-top colors (red, white, gray, and blue) and varied the camera’s heights (150cm, 177cm, and 185cm) using ARMAR-6’s torso as demonstrated in Fig. 3. Additionally, we recorded data under three different pitch angles (10 degrees, 37 degrees, and 49 degrees down) and six different lighting conditions as shown in Fig. 4. To avoid similar and repetitive frames, we limited our recording to 5 $fps$. To the best of our knowledge, this is the first of its kind dataset that covers this amount of different robots’ fields of view.
@@ III-C Annotation Pipeline
Figure 5: Our proposed Annotation Pipeline. The pipeline begins with inputting 3D meshes of dataset objects, which BlenderProc2 processes to generate synthetic data with 2D bounding boxes. This annotated 2D data is utilized to train a YOLOv5 2D object detector. Subsequently, real-world recorded data is fed into the trained model, and the output undergoes manual inspection for correct and incorrect labeling. The correctly labeled images are employed for model refinement, which is then validated on the incorrectly labeled ones. This iterative process continues until all images are accurately labeled. The images with correct labels are then passed to Segment Anything (SAM) to produce masks. Finally, the images, along with the 2D labels and 3D meshes, are input into MegaPose to generate 6D poses for detected objects. Manual inspection of poses is conducted through contour and mesh overlay images, and corrected annotations are used to fine-tune MegaPose iteratively until the entire dataset is accurately annotated.
Annotating objects with their ground truth 6D poses is a labor-intensive and time-consuming task. To streamline this process, we propose a semi-automated annotation pipeline. This pipeline generates three types of annotations: 2D object bounding boxes, 2D segmentation masks, and 6D poses.
@@ III-D Comparison to Existing Datasets
When compared to currently available datasets, the KITchen dataset stands out in several key aspects. With a diverse collection of 111 objects, our dataset offers a significantly wider range than the average number of objects found in existing datasets, surpassing the average by a factor of four. This expansive variety is crucial for training robust pose estimation models capable of handling a multitude of real-world scenarios. Moreover, the KITchen dataset offers a total of 205K RGBD images. This surpasses the average number of annotated images in existing datasets by over threefold, providing more data for training and evaluation purposes. Furthermore, our dataset has a remarkably larger number of annotated objects per image compared to the existing datasets with an unprecedented number of objects reaching 50 per image. This exceeds any available dataset by a significant margin, enabling more comprehensive analysis and training of instance-level 6D pose estimation models. Additionally, the KITchen dataset is unique in its capture methodology. It is the only dataset to have been recorded using the field of view of a humanoid robot with adjustable heights, camera angles, and lighting conditions. Unlike existing datasets that predominantly focus on tabletop scenes, our dataset features challenging locations within kitchen environments including refrigerators, ovens, sinks, higher shelves, microwaves, and dishwashers, offering a broader scope of real-world scenarios for pose estimation research. An overview of the dataset comparison is given in Tab. I.
@@ IV-B Datasets
Our benchmark leverages the KITchen dataset introduced in Sec. III. Notably, this dataset stands out as the first of its kind, captured from the perspective of a humanoid robot, and encompasses varying heights and pitch angles, making it more suited to cover robotic mobile manipulation scenarios in kitchen environments. We split the dataset to train/val/test sets with a 70/20/10 ratio.
@@ V Conclusion
We introduce KITchen, a novel object 6D pose estimation benchmark tailored to tackle this task within challenging kitchen environments using only monocular vision from robots’ FOV, with a specific emphasis on real-time performance. To serve this benchmark, we recorded the first-of-its-kind large-scale real-world dataset, captured from the perspective of a humanoid robot, featuring multi-objects in structured cluttered scenes in two distinct kitchen environments with diverse lighting conditions. Lastly, we proposed a semi-automated annotation pipeline aimed at streamlining the annotation of such datasets while minimizing manual human effort. We envision our benchmark as a bridge between robotics and the computer vision fields, fostering the development of innovative approaches to solve the 6D pose problem on resource-constrained platforms while also prioritizing real-time applicability.

### v2

@@ Abstract
Despite the recent progress on 6D object pose estimation methods for robotic grasping, a substantial performance gap persists between the capabilities of these methods on existing datasets and their efficacy in real-world grasping and mobile manipulation tasks, particularly when robots rely solely on their monocular egocentric field of view (FOV). Existing real-world datasets primarily focus on table-top grasping scenarios, where a robot arm is placed in a fixed position and the objects are centralized within the FOV of fixed external camera(s). Assessing performance on such datasets may not accurately reflect the challenges encountered in everyday grasping and mobile manipulation tasks within kitchen environments such as retrieving objects from higher shelves, sinks, dishwashers, ovens, refrigerators, or microwaves. To address this gap, we present KITchen, a novel benchmark designed specifically for estimating the 6D poses of objects located in diverse positions within kitchen settings. For this purpose, we recorded a comprehensive dataset comprising around 205k real-world RGBD images for 111 kitchen objects captured in two distinct kitchens, utilizing a humanoid robot with its egocentric perspectives. Subsequently, we developed a semi-automated annotation pipeline, to streamline the labeling process of such datasets, resulting in the generation of 2D object labels, 2D object segmentation masks, and 6D object poses with minimal human effort. The benchmark, the dataset, and the annotation pipeline will be publicly available at https://kitchen-dataset.github.io/KITchen.
@@ I Introduction
Figure 2: The humanoid robot ARMAR-6, leveraged for its adjustable torso height and various camera angles provided by its adjustable roll-yaw neck, to enrich our dataset.
To overcome the limitations of the current 6D pose estimation methods, we introduce KITchen, the first-of-its-kind large-scale real-world dataset recorded using the humanoid robot ARMAR-6 [2] as shown in Fig. 2, which has adjustable height and roll-yaw neck, in 2 different kitchen environments covering 111 kitchen objects from the robots’ egocentric perspective to cover the objects in the challenging kitchens’ locations as shown in Fig. 1. KITchen offers 2D bounding boxes, object segmentation, and 6D poses annotated with a semi-automated annotation pipeline to minimize the need for manual labeling.
The main contributions of our work are: (i) we introduce a large real-world annotated RGBD dataset for 111 objects with their 2D bounding boxes, segmentation masks, and 6D poses. (ii) we propose a semi-automated annotation pipeline to annotate the objects in the dataset to facilitate the creation of more real-world datasets and make it publicly available to other researchers to create such large-scale datasets. (iii) we introduce a new benchmark and competition, where the focus is to solve the object 6D pose estimation problem depending solely on the monocular FOV of robots and limiting the submissions to approaches that offer at least $5fps$ to encourage further work on this problem while taking into consideration real-time applicability.
@@ II Related Work
Dataset Objects Images Annotated Objects/Image Multi-object Multi-instance Mobile Robot’s FOV
@@
TABLE I: Overview of available datasets for instance-level 6D pose estimation
@@ II-A Objects Datasets
GraspNet-1Billion [18] contains around 97.3K RGBD images for 88 objects recorded with 2 different cameras for table-top grasping scenario with one robot arm. KIT object models database [29] was originally introduced in 2012 and offers 3D CAD models for more than 100 diverse objects, the majority of which are kitchen-related groceries. However, it only offers very few images for each object, which makes it hard to use this dataset for 6D pose estimation with the current state-of-the-art (SOTA) data-driven 6D pose estimation approaches. KIT bimanual manipulation dataset [31] provides rich data for learning models of bimanual manipulation tasks from human demonstrations. It includes accurate whole-body motion data, hand configurations, and 6D object poses captured using various sensors. The dataset features 12 bimanual actions for 21 kitchen-related objects.
An overview of available datasets for instance-level 6D pose estimation is given in  Table I. The overview highlights key metrics including the number of covered objects in the dataset, total image count, number of annotated objects per image, presence of multi-object setups, availability of multiple instances of the same objects, and whether the dataset was captured using a mobile robot’s field of view.
@@ III-A Dataset’s Objects
We aim to create a large-scale real-world dataset that covers objects that are commonly used in kitchen environments. Although some of the existing object datasets already offer objects that are commonly used in kitchens, they lack enough diverse RGBD annotated images to train on [50] or no annotated RGBD at all [6, 31, 29, 35]. Therefore, we decided to reuse the already available kitchen-related objects from these datasets and provide a large real-world RGBD annotated dataset for them to facilitate research on 6D pose estimation for kitchen objects. These objects vary from toy vegetables and fruits from [6] to kitchen tools such as knives, spoons, cups, mugs, bowls, cutting board, egg whisk, frying pan, plate, etc. from [6, 50, 31, 29, 35] to kitchen groceries objects from [6, 29].
@@ III-B Dataset Recording
Figure 4: Diverse robot and camera heights realized through different torso positions of ARMAR-6. The images display heights of 145cm, 177cm, and 185cm from left to right, illustrating the varied perspectives captured in the datasets and the different placements of objects relative to the robot’s field of view.
We recorded the dataset using our humanoid robot ARMAR-6 [3] inside two distinct kitchen environments as seen in Fig. 3 the first kitchen, referred to as the Main Kitchen, includes typical kitchen appliances such as a fridge, counter with drawers, table, sink, microwave, dishwasher, and oven. The second kitchen, named Mobile Kitchen, features a counter with drawers, sink, dishwasher, fridge, and three tables. To enhance diversity, we utilized four different table-top colors (red, white, gray, and blue) and varied the camera’s heights (150cm, 177cm, and 185cm) using ARMAR-6’s torso as shown in Fig. 4. Additionally, we recorded data under three different pitch angles (10 degrees, 37 degrees, and 49 degrees down) and six different lighting conditions as shown in Fig. 5. We shuffled the objects with each change of lighting, camera’s height, or camera’s angle to enrich the diversity of the recorded scenes. To avoid similar and repetitive frames, we limited our recording to 5 $fps$. To the best of our knowledge, this is the first of its kind dataset that covers this amount of different robots’ fields of view.
@@ III-C Annotation Pipeline
Annotating objects with their ground truth 6D poses is a labor-intensive and time-consuming task. Although some of the recently published datasets attempted to semi-automate the annotation process. For instance, GraspNet-1Billion’s approach [18] relies on manually annotating the first frame of each scene, then leveraging recorded camera poses to calculate the objects’ poses in the following frames. However, this method was not optimal for our dataset, as KITchen has many more diverse scenes per kitchen compared to the simple setup used in GraspNet-1Billion, resulting in significantly more effort required for manual annotation. Another attempt to semi-automate the annotation process was presented by HANDAL [19], but their approach assumes a single object in the scene, making it unsuitable for our dataset, which contains multiple objects $10-50$ objects in each scene. To overcome the limitations of existing approaches and streamline the annotation process, we propose a semi-automated annotation pipeline. This pipeline generates three types of annotations: 2D object bounding boxes, 2D segmentation masks, and 6D poses, see 6.
Figure 6: Our proposed annotation pipeline. The pipeline starts 3D meshes of dataset objects as input, which are processed by BlenderProc2 to generate synthetic data with 2D bounding boxes. This annotated 2D data is used to train a YOLOv5 2D object detector. Subsequently, real-world recorded data is fed into the trained model, and the output is manually inspected for correct and incorrect labeling. The correctly labeled images are used for model refinement, which is then validated on the incorrectly labeled images. This iterative process continues until all images are correctly labeled. The correctly labeled images are then passed to Segment Anything (SAM) to generate masks. Finally, the images, along with the 2D labels and 3D meshes, are fed into MegaPose to generate 6D poses for detected objects. Manual inspection of poses is performed using contour and mesh overlay images, and corrected annotations are used to iteratively fine-tune MegaPose until the entire dataset is accurately annotated.
@@ III-D Comparison to Existing Datasets
When compared to currently available datasets, the KITchen dataset stands out in several key aspects. With a diverse collection of 111 objects, our dataset offers a significantly wider range than the average number of objects found in existing datasets, surpassing the average by a factor of four. This expansive variety is crucial for training robust pose estimation models capable of handling a multitude of real-world scenarios. Moreover, the KITchen dataset offers a total of 205K RGBD images. This surpasses the average number of annotated images in existing datasets covered in Table I by over threefold, providing more data for training and evaluation purposes. Furthermore, our dataset has a remarkably larger number of annotated objects per image compared to the existing datasets with an unprecedented number of objects reaching 50 per image. This exceeds any available dataset by a significant margin, enabling more comprehensive analysis and training of instance-level 6D pose estimation models. Additionally, the KITchen dataset is unique in its capture methodology. It is the only dataset to have been recorded using the field of view of a humanoid robot with adjustable heights, camera angles, and lighting conditions. Unlike existing datasets that predominantly focus on tabletop scenes, our dataset features challenging locations within kitchen environments including refrigerators, ovens, sinks, higher shelves, microwaves, and dishwashers, offering a broader scope of real-world scenarios for pose estimation research. An overview of the dataset comparison is given in Table I.
@@ IV-B Datasets
Our benchmark leverages the KITchen dataset introduced in Sec. III. Notably, this dataset stands out as the first of its kind, captured from the perspective of a humanoid robot, and encompasses varying heights and pitch angles, making it more suited to cover robotic mobile manipulation scenarios in kitchen environments. We split the dataset to training/validation/test sets with a 70/20/10 ratio.
@@ V Conclusion
We introduce KITchen, a novel object 6D pose estimation benchmark tailored to tackle this task within challenging kitchen environments using only monocular vision from robots’ FOV, with a specific emphasis on real-time performance. To serve this benchmark, we recorded a large-scale real-world dataset, captured from different perspectives of a humanoid robot, featuring multi-objects in structured cluttered scenes in two distinct kitchen environments with diverse lighting conditions. Lastly, we proposed a semi-automated annotation pipeline aimed at streamlining the annotation of such datasets while minimizing manual human effort. We envision our benchmark to promote the development of novel approaches to solve the 6D pose problem on resource-constrained platforms, with an emphasis on real-time and real-world applicability.